OpenStudy (anonymous):

what is the period of the function y=2 cos pi x

OpenStudy (anonymous):

period of \[\cos(bx)\] is \[\frac{2\pi}{b}\] in your example \(b=\pi\)

OpenStudy (anonymous):

So the period is pi?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

the period is not \(b\) it is \(\frac{2\pi}{b}\)

OpenStudy (anonymous):

So it's just 2? since b=pi

OpenStudy (anonymous):

yes 2

OpenStudy (anonymous):

Awesome, thanks. Can you help me with some more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what are the graphs of y = sin x and y = csc x in the interval from –2pi to 2pi

OpenStudy (anonymous):

@satellite73